# Quantum mechanics provides the physical basis of teleological evolutions

### v2

@@ 1 Introduction
Let us briefly examine the very long history of the teleological notion—attraction by a final cause. Plato and Aristotle used it to explain the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution. With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy. It was eventually dismissed by science for the alleged absence of a physical basis. All this was before the advent of quantum mechanics. Today, the situation of the teleological notion is somewhat complex. To the best of our knowledge, outside the present work, there are no physical researches on it. Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science. This is because, in the first place, quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since showing this will be surprisingly simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
@@
By the way, here and in the following by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible. Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem).
@@
As a matter of fact, we have provided two redundant demonstrations of this rule, what should not be despised given its destabilizing character. Let us outline them.
@@
By the way, we will have to use a quantum description of the quantum algorithms that is not their customary quantum description. The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement. We complete the customary quantum description by adding to the description of the process of solving the problem that of the upstream process of setting the problem. The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow. Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic. As we shall see in the next section, the classical logic description of the quantum speedup is indeed the advanced knowledge rule.
@@
Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete. Indeed, it is simply completed by time-symmetrizing it. We shall see that the time-symmetrization of the quantum description yields the quantum superposition of the causal loops involved in the advanced knowledge rule. In other words, the time-symmetrization of the quantum description of the quantum algorithm generates the physical correspondent of the advanced knowledge rule.
@@
The second demonstration of the advanced knowledge rule simply comes from the fact that this time-symmetrization of the quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far. As we shall see, it is to satisfy an atemporal symmetry implicit in the quantum description. Of course, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—it is a demonstration of the rule.
@@
Of course, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
@@
A further step, important for what concerns the possible teleological character of the evolutions of the living, is that all the above generalizes to quantum correlation. For simplicity, we limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty. We can then say that the quantum algorithm builds up one-to-one correlation between two initially unrelated observables (the problem-setting and the solution of the problem) without ever changing the eigenvalue of the observable selected by the initial measurement (the problem-setting). We shall show that any building up of quantum correlation of this kind (not only that in quantum algorithms) can occur with quantum speedup and therefore be teleological in character. Note that the “can be” means that there is certainly room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation. By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
With this, we can go to the evolutions of the living for which the teleological notion was originally conceived.
@@
Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living. The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, interestingly, the very context we are dealing with suggests a way out. The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11]. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such, it can occur with quantum speedup and thus be teleological in character.
@@
The Fine-tuned universe version of the Anthropic Principle is about the very relevant fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
@@
Under the quantum cosmological assumption, this version of the Principle can be formalized as follows. We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
Now, always under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental physical laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental physical constants, and each with its own value of the latter. Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”. The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life. Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life; in the other universes, where this value is different, there is absence of life (theory says that, the absence of life in the other universes today is already definitive).
@@
For what we have said about the generalization of the quantum speedup from quantum algorithms to the building up of quantum correlation, the evolution of the universe toward the value of life can occur with quantum speedup and, therefore, be teleological in character. The “can occur” is in the sense that there is room for its occurring with quantum speedup.
@@
Of course, in the universe we are in, the possible quantum speedup and the consequent teleological character is for the evolution of the universe toward life. We will see that this possibility would be inherited by the evolution of life on Earth, in principle, the trace over all the rest of the evolution of the universe toward life. Then, for consistency, it would be inherited by the evolutions of species.
@@
Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered as plausible as the quantum cosmological assumption. However, note that all this is quite different from the present widespread belief that the notion of teleological evolution of the living is an absurdity.
@@
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe obviously holds only in principle and does not give what the classical description gives. However, this might pinpoint a direction of the research. At the theoretical level, it could be interesting to study how this “tracing over all the rest” conciliates a holistic and teleological view in the longer term with a mechanistic and reductionist view in the shorter term. This might even be a further perspective under which to explore the unification between classical and quantum mechanics.
@@ 2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality. Since this is a way of thinking quantum computation radically different from the usual one, we provide it first at a conceptual level. All formalizations are postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. This said, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
@@ 2.1 On the explanation of the quantum computational speedup
For what concerns the current mainstream explanation of the quantum speedup, let us say, it is the case of crying out that the emperor is naked. This emphasis is to underline the absence of any recognition of the fact that the current view of quantum computation has a shortcoming.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against the deeply rooted de facto ban on retrocausality. In this work, we should have demonstrated it, moreover, in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the unconventional character of the rule.
@@ 2.2 First demonstration of the advanced knoledge rule
We have already given in the Inroduction the quantum description of quantum algorithms extended to the process of solving the problem. An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow. For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
Now, it should also be noted that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. To present ends, we should also relativize it with respect to the problem-solver, who computes the solution of the problem. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her. Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easily obtained. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of the unitary evolution.
This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
@@
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [12]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [13] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference. By the way, note that the notion of searching for something and finding it is most universal and this universality is completely represented by Grover algorithm.
We show that the classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule. We begin with the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
@@
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings); the fewest is the number needed in the optimal logical case benefiting of half of the information about the solution. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting from this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm.

### v3

@@ 1 Introduction
Let us briefly retrace the two and a half millennia history of the teleological notion: attraction exercised by a final cause. Plato and Aristotle applied it the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution. With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy. It was eventually dismissed by science (it was classical science) for the alleged absence of a physical basis, even ascribed, for its finalistic character, to the category of beliefs. All this was before the advent of quantum mechanics, but has remained almost unchanged with it. Today, the situation of the teleological notion is somewhat complex. However, to the best of our knowledge, outside the present work, in physics, there are no researches on it. Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science. This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since showing this will be simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
@@
By the way, here and in the following, by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible. Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem). We will see further on that the physical definition of teleological evolution must be that of an evolution between an initial and a final measurement outcome.
@@
As a matter of fact, we will provide two redundant demonstrations of this rule, which should not be despised given its destabilizing character. Let us outline them.
@@
By the way, we will have to use a quantum description of the quantum algorithms that is not the customary one. The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement.
We complete this customary quantum description by adding, to the description of the process of solving the problem, that of the upstream process of setting the problem. The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow. Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic.
Having said this, the outline of the first demonstration can be very short. As we shall see in the next section, Under the present more complete quantum description of the quantum algorithm, the classical logic description of the quantum speedup turns out to be indeed the advanced knowledge rule.
@@
Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete. Indeed, it is simply completed by time-symmetrizing it. We shall see that the time-symmetrization of the more complete quantum description of the quantum algorithm generates the quantum superposition of the causal loops involved in the advanced knowledge rule.
@@
The second demonstration of the advanced knowledge rule simply comes from the fact that the time-symmetrization of the more complete quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far. As we shall see, it is to satisfy an atemporal symmetry implicit in this description. As we have seen, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—we have thus a demonstration of the rule.
@@
Naturally, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
A further step, important for the possibility of extending the teleological character of quantum algorithms to the evolutions of the living, is the following generalization.
@@
For simplicity, let us limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty. We can then say that the quantum algorithm unitarily builds up one-to-one correlation between two initially unrelated observables—the problem-setting and the solution of the problem—without ever changing the eigenvalue of the observable selected by the initial measurement—the problem-setting.
We shall show that any building up of quantum correlation of this kind, also outside quantum computation, can occur with quantum speedup and therefore be teleological in character. Note that the “can occur” means that there is certainly the room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation. By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
We go now to the evolutions of the living for which the teleological notion was conceived.
@@
Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living. The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, the very context we are dealing with suggests a way out. The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11]. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) could acquire the form of a quantum algorithm and, as such, it could occur with quantum speedup and thus be teleological in character.
@@
The Fine-tuned universe version of the Anthropic Principle is about the extraordinary theoretical fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
Under the quantum cosmological assumption, this version of the Principle can be formalized as follows.
@@
We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
Of course, at the beginning of time, these two observables must be unrelated: no matter the value of the fundamental constant, the value of life is “absence of life”. At the time of today, the unitary evolution of the universe toward the value of life has built up one-to-one quantum correlation between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value.
@@
We can therefore say that the evolution of the universe toward the value of life can occur with quantum speedup and thus be teleological in character; the “can occur” is in the sense that there is certainly the room for its occurring with quantum speedup. Of course, whether the quantum speedup is achieved depends on the character of the unitary evolution. For what was said about quantum random search algorithms, we might think that, if the search for the value of life by the unitary evolution of the universe is random, it should occur with a quantum speedup.
@@
In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; of course, all the above also applies to the latter evolution. We shall see that, as a consequence, the evolution of life on Earth, in principle the trace over all the rest of the evolution of the universe toward life, would inherit the possible teleological character of the evolution of the universe toward life. For consistency, the same would apply to the evolution of species.
@@
Of course, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption. Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity. At least, it should cast a doubt on the present position of science.
@@
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe, obviously, holds only in principle: in practice, we cannot have such descriptions. However, this contrast between the quantum and the classical view might pinpoint a direction of the research. At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system in the longer term with a classical character in the shorter term. This might even be a further perspective under which to study the unification between quantum and classical mechanics.
@@ 2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality. Since this is a way of thinking quantum computation that is radically different from the usual one, we provide it first at a conceptual level. All formalizations are postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
@@ 2.1 On the explanation of the quantum computational speedup
Summing up, in our judgment, for what concerns the current mainstream explanation of the quantum speedup, it is the case of saying that the emperor is naked.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it has not been demonstrated; second, it goes against the de facto ban on retrocausality. In this work, we should have demonstrated it, moreover, in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent demonstrations, which is not to be despised because of the destabilizing character of the rule.
@@ 2.2 First demonstration of the advanced knoledge rule
We have already provided in the Introduction the quantum description of quantum algorithms extended to the process of solving the problem. An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow. For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
Now, we should note that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. To present ends, we should relativize it with respect to the problem-solver who, for the fact of computing the solution of the problem, is part of the process. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the completed quantum description of the quantum algorithm with respect to her. Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would tell her the solution of the problem before she begins her problem-solving action. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easy. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Section (2) of the Appendix.
@@
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [12]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [13] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference. By the way, note that the notion of searching for something and finding it is most universal and that this universality is completely represented by Grover algorithm.
We show that the classical logic description of the quantum speedup delivered by Grover algorithm very directly implies the advanced knowledge rule. We begin with the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
@@
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future (that the ball is in two rather than four drawers) and can use this knowledge to produce the solution by opening just one drawer. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting from this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm. This is, of course, the advanced knowledge rule for the full fledged Grover algorithm.